### 16 Mar 2024 v1

# Zero Lyapunov exponents in transitive skew-products of Iterated Function systems

Abstract.
We study the class of transitive skew-products associated with iterated function systems of circle diffeomorphisms. We can approximate any transitive skew-product by maps in this class that have a robustly zero Lyapunov exponent. In particular, we prove the existence of non-hyperbolic ergodic measures for an open and dense subset of transitive skew-products. Moreover, these measures have full support and are the $*{}^{*}$ limit of periodic measures.
1. Introduction
The theory of uniformly hyperbolic dynamical systems was initiated in the 1960s by Smale [Sma67] and Anosov [Ano67], providing a detailed description of a large class of systems, often with very complex evolution. Since the 1970s, with the works of Abraham and Smale [AS69], and Newhouse [New70], it has been known that uniformly hyperbolic systems do not form a dense set of dynamical systems. In response to the absence of density in these spaces, weaker concepts emerged, such as the non-uniformly hyperbolic systems introduced by Pesin [Pes77] in 1977. These systems are characterized by having hyperbolic ergodic measures, that is, with non-zero Lyapunov exponents. It took a few decades to demonstrate that non-uniformly hyperbolic systems are also not dense. Namely, Kleptsyn and Nalsky [KN07] constructed in 2007 a dynamical system on the torus $\mathbb{T}^{3}$ with a robustly zero Lyapunov exponent. Later, in 2016, Bochi, Bonatti, and Díaz [BBD16] presented new open sets of diffeomorphisms in higher-dimensional manifolds with non-hyperbolic ergodic measures, that is with zero exponents. The result shown by Kleptsyn and Nalsky is based on a toy model built a few years earlier by the Russian school of Ilyashenko [${}^{+}$05]. In this paper, we will show that the existence of non-hyperbolic ergodic measures is abundant in the class of transitive skew-products associated with iterated function systems where the toy models in [${}^{+}$05] were constructed.
An iterated function system (IFS) on $\mathbb{S}^{1}$ is a family of functions =$\mathscr{F}=\{f_{1},\dots,f_{k}\}$ from the circle $\mathbb{S}^{1}$ to itself that can be applied (composed) successively in any order. Therefore, such compositions are elements of the semigroup $\langle\mathscr{F}\rangle^{+}$ associated with these transformations. Moreover, the sequence of compositions can be seen as the fiber of the iterations of the locally constant skew-product
$F:\Sigma_{k}\times\mathbb{S}^{1}\to\Sigma_{k}\times\mathbb{S}^{1},\qquad F^{n}(\omega,x)=(\sigma^{n}(\omega),f^{n}_{\omega}(x))$ (1)
where at the base we have the shift $\sigma:\Sigma_{k}\to\Sigma_{k}$ and
$f^{n}_{\omega}\stackrel{{\scriptstyle\scriptscriptstyle\rm def}}{{=}}f_{\omega_{n-1}}\circ\dots\circ f_{\omega_{0}}\quad\text{for}\ \ \omega=(\omega_{i})_{i\geq 0}\in\Sigma_{k}\stackrel{{\scriptstyle\scriptscriptstyle\rm def}}{{=}}\{1,\dots,k\}^{\mathbb{N}}\quad\text{and}\quad n\geq 1.$
To emphasize the role of the fiber maps and the product structure, we will write $F=\sigma\ltimes\mathscr{F}$.
Given an ergodic $F$-invariant measure $\mu$, Birkhoff’s Ergodic Theorem guarantees the existence of the limit
$\lambda(\omega,x)=\lim_{n\to\infty}\frac{1}{n}\log|(f^{n}_{\omega})^{\prime}(x)|\quad\text{for $\mu$-a.e.~{}$(\omega,x)\in\Sigma_{k}\times\mathbb{S}^{1}$}.$
In fact
$\lambda(\mu)\stackrel{{\scriptstyle\scriptscriptstyle\rm def}}{{=}}\int\log|(f_{\omega_{0}})^{\prime}(x)|\,d\mu=\lambda(\omega,x)\quad\text{for $\mu$-a.e $(\omega,x)\in\Sigma_{k}\times\mathbb{S}^{1}$}.$
The value $\lambda(\mu)$ is called the Lyapunov exponent along the fiber of the skew product $F$ with respect to the measure $\mu$.
For $r\geq 1$ and $k\geq 2$, we denote by $\operatorname{\mathrm{IFS}}_{k}^{r}(\mathbb{S}^{1})$ the set of families =$\mathscr{F}=\{f_{1},\dots,f_{k}\}$ with $f_{i}$ belongs to the space $\mathrm{Diff}^{r}_{+}(\mathbb{S}^{1})$ of orientation-preserving diffeomorphisms of the circle equipped with the $C^{r}$-topology. Moreover, we say that =$\mathscr{F}=\{f_{1},\dots,f_{k}\}$ and $\mathscr{G}=\{g_{1},\dots,g_{k}\}$ are $C^{r}$-close if, by reordering the elements families if necessary, $f_{i}$ and $g_{i}$ are $C^{r}$-close for every $i=1,\dots,k$. We define $\mathrm{S}_{k}^{r}(\mathbb{S}^{1})$ as the set of skew-products $F=\sigma\ltimes\mathscr{F}$, where $\mathscr{F}\in\operatorname{\mathrm{IFS}}_{k}^{r}(\mathbb{S}^{1})$. We endow $\mathrm{S}_{k}^{r}(\mathbb{S}^{1})$ with a topology, stating that $F=\sigma\ltimes\mathscr{F}$ is close to $G=\sigma\ltimes\mathscr{G}$ if $\mathscr{F}$ and $\mathscr{G}$ are $C^{r}$-close.
Gorodetski et al. [${}^{+}$05] constructed an open set $\mathcal{U}$ in $\mathrm{S}_{2}^{1}(\mathbb{S}^{1})$ such that each ${F}\in\mathcal{U}$ has an ergodic invariant measure $\mu$ with $\lambda(\mu)=0$. Originally in [${}^{+}$05], it is shown that this open set can be constructed as a $C^{r}$-neighborhood of a skew-product $G=\sigma\ltimes\mathscr{G}$ with $\mathscr{G}=\{g_{1},g_{2}\}$ where $g_{1}$ is an irrational rotation and $g_{2}$ has a pair of hyperbolic periodic points with additional conditions on the eigenvalues. Later, Bochi et al. [BBD13] generalized this result for IFSs on compact manifolds of higher dimension. In fact, they constructed open sets of IFSs in the $C^{r}$ topology for $r\geq 2$ where the associated skew-products have an ergodic invariant measure, limit in the weak-star topology of periodic measures, with full support and where all Lyapunov exponents along the fibers are zero.
A skew-product $F\in\mathrm{S}_{k}^{r}(\mathbb{S}^{1})$ is said to be transitive if there has a dense orbit. We denote by $\mathrm{TS}_{k}^{r}(\mathbb{S}^{1})$ the subset of $\mathrm{S}_{k}^{r}(\mathbb{S}^{1})$ formed by transitive skew-products. Note that the open set of skew-products constructed in [${}^{+}$05] belongs to this set. The following main result shows that any transitive skew-product can be approximated by an open (in $\mathrm{S}^{r}_{k}(\mathbb{S}^{1})$) set of transitive skew-products having a non-hyperbolic ergodic measure with full support. This concludes affirmatively the conjecture [DGR17, Conj. 2] in our setting.
Theorem A.
For $k\geq 2$ and $1\leq r\leq\infty$, there exists an open and dense subset $\mathcal{R}$ of $\mathrm{TS}_{k}^{r}(\mathbb{S}^{1})$ such that for every $F\in\mathcal{R}$ there is an ergodic $F$-invariant probability measure $\mu$ with full support whose Lyapunov exponent $\lambda(\mu)$ equal to zero. The probability measure $\mu$ is a limit in the $*{}^{*}$ topology of sequences of invariant measures supported on periodic points.
The $C^{1}$-density of maps with non-hyperbolic ergodic invariant measures, as in the above theorem, was stated in [DGR17, Cor. 2 and 8.10] on the set of all locally constant [^1] Although the result in [DGR17] is stated for Hölder skew-products, the approximation by locally constant skew-products in cylinders proposed is not clear. skew-products with $C^{1}$-fiber maps that are robustly transitive and have periodic points of different indices. This result is a corollary of Proposition 8.9 in [DGR17], which is not fully proved. The authors observe that the proof of a similar result in [BDU02] for partially hyperbolic diffeomorphisms of compact manifolds can be translated mutatis mutandis to obtain such proposition in the locally constant skew-product setting. The results in [BDU02] show the density of the strong unstable (strong stable) foliation, and the authors in [DGR17] translate this property to the locally constant skew-product setting as the forward (backward) minimality of the underlying IFS. However, this is not entirely true, as the results in [BC23, Cor. 6.8] have shown. Nevertheless, according to [BC23, Thm. B], it follows that the set of skew-products having both strong stable and strong unstable foliations also contains an open and dense subset of $\mathrm{TS}^{r}_{k}(\mathbb{S}^{1})$. Consequently, $\mathcal{R}$ in Theorem A can also possess this property.
To prove Theorem A we need to the next slightly generalization of [${}^{+}$05, Thm. 2] for IFSs having more than two generators. First, we recall some definitions. A finite family $\mathscr{F}$ of circle diffeomorphisms is minimal if for every $x\in\mathbb{S}^{1}$ the orbit $\mathcal{O}(x)=\{f(x):f\in\langle\mathscr{F}\rangle^{+}\}$ is a dense set in $\mathbb{S}^{1}$. Also, we say that $\mathscr{F}$ is backward (resp. forward) expanding if for every $x\in\mathbb{S}^{1}$, there exists $f$ in $\langle\mathscr{F}\rangle^{+}$ (resp. $\langle\mathscr{F}^{-1}\rangle^{+}$, $\mathscr{F}^{-1}=\{g^{-1}:g\in\mathscr{F}\}$) such that $|f^{\prime}(x)|>1$.
Theorem B.
For any $k\geq 2$, consider the skew-product $F=\sigma\ltimes\mathscr{F}$ where =$\mathscr{F}=\{f_{1},\dots,f_{k}\}$ is a family of circle $C^{1}$-difeomorphisms such that
  1. (i)
    $\mathscr{F}$ is minimal;
  2. (ii)
    $\mathscr{F}$ is backward expanding;
  3. (iii)
    there exists $f\in\langle\mathscr{F}\rangle^{+}$ with a hyperbolic attracting periodic point.
Then $F$ has an ergodic invariant measure $\mu$ with full support, limit of invariant measures supported on periodic points and $\lambda(\mu)=0$.
The property of full support of the measure was not directly proven in [${}^{+}$05]. This property was obtained later in [BBD13] by spreading the sequence of periodic points in the ambient space and using results from [BDG10]. We replicate these ideas here to demonstrate the above result.
It is not difficult to see that properties (ii) and (iii) in the preceding theorem are open in the $C^{1}$ topology. This means that every $C^{1}$-close IFS still satisfies these properties. However, property (i) is not generally open. Despite this, it was established in [${}^{+}$17] that if $\mathscr{F}$ is both minimal and forward expanding, then it is also minimal in a $C^{1}$-robust sense. Since the minimality of $\mathscr{F}$ implies the transitivity of $F=\sigma\ltimes\mathscr{F}$, c.f. [BC23, Remark 4.4 and Theorem 6.1], we arrive at the following remark:
Remark 1.1.
The set of skew-products $F=\sigma\ltimes\mathscr{F}\in S^{r}_{k}(\mathbb{S}^{1})$ such that $\mathscr{F}$ satisfies (i)–(iii) and
  • (iv)
    $\mathscr{F}$ is forward expanding
is an open subset of $\mathrm{TS}^{r}_{k}(\mathbb{S}^{1})$ for every $r\geq 1$.
Remark 1.2.
An open set $B$ is called globalized blending region for $\mathscr{F}$ if there exist maps $h_{1},\dots,h_{m},T_{1},\dots,T_{s},S_{1}\dots,S_{t}\in\langle\mathscr{F}\rangle^{+}$ and an open set $D\subset\mathbb{S}^{1}$ such that $\overline{B}\subset D$ and
  1. (1)
    $\overline{B}\subset h_{1}(B)\cup\dots\cup h_{m}(B)$,
  2. (2)
    $h_{i}:\overline{D}\to D$ is a contracting map for $i=1,\dots,m$,
  3. (3)
    $\mathbb{S}^{1}=T_{1}(B)\cup\dots\cup T_{s}(B)=S_{1}^{-1}(B)\cup\dots\cup S_{t}^{-1}(B)$.
According to [${}^{+}$17, Sec. 6], a skew-product $F=\sigma\ltimes\mathscr{F}$, where both $\mathscr{F}$ and $\mathscr{F}^{-1}$ have a globalized blending region, satisfies (i)–(iv). In particular, Theorem B applies. This can be compared with [DGR17, Prop. 8.3].
In the following section, §2, we will prove Theorem B by using some preliminary results from [${}^{+}$05, DG09, BDG10, BBD13]. Afterward, in §3, guided by Theorem B and Remark 1.1, the fundamental idea behind the proof of Theorem A is to establish that, up to conjugation, any transitive locally constant skew-product can be approximated by one satisfying conditions (i) through (iv). These properties essentially arise from constructing a pair of globalized blending regions for $\mathscr{F}$ and $\mathscr{F}^{-1}$. This construction is found within the proof of the results borrowed from [KKO18, BC23] to prove Theorem A.
2. Proof of Theorem B
Fix $k\geq 1$ and consider a family =$\mathscr{F}=\{f_{1},\dots,f_{k}\}$ of circle diffeomorphisms.
2.1. Preliminares
The next proposition proved in [${}^{+}$05, Lem. 1] is key to understanding the Lyapunov exponent approximation method which will be developed in the next subsection.
Proposition 2.1.
Let $\{\mu_{n}\}_{n\in\mathbb{N}}$ and $\mu$ be invariant ergodic probability measures for the skew-product $F=\sigma\ltimes\mathscr{F}$ given in (1). Then, if $\mu_{n}\to\mu$ as $n\to\infty$, it follows that $\lambda(\mu_{n})\to\lambda(\mu)$ as $n\to\infty.$
The approach suggested in [${}^{+}$05] and and [BDG10] allows to construct non-hyperbolic ergodic invariant measures using the previous proposition as limits of measures supported on special sequences of periodic points. Recall that given an $n$-periodic orbit $X=\left\{G(x_{0}),G^{2}(x_{0}),\dots,G^{n-1}(x_{0}),G^{n}(x_{0})=x_{0}\right\}$ of a continuous map $G$ of a compact metric space, atomic measure uniformly distributed on $X$ is given by
$\mu_{X}=\frac{1}{n}\sum_{i=0}^{n-1}{\delta_{{G}^{i}(x_{0})}}$
where ${\delta}_{y}$ is the Dirac measure at point $y$. This measure is $G$-invariant and ergodic.
Definition 2.2.
Let $X$ and $Y$ be periodic orbits of $G$. Given $\gamma>0$ and $\aleph>0$, we say that $Y$ is a $(\gamma,\aleph)$-good approximation of $X$ if the following conditions are met:
  1. (i)
    There exists a subset $\Gamma$ of $Y$ and a projection $\rho:\Gamma\to X$ such that $d(G^{j}(y),G^{j}(\rho(y)))<\gamma$ for all $y\in\Gamma$ and all $j=0,1,\dots,P-1$. Here, $P$ is the period of the orbit $X$;
  2. (ii)
    #$\frac{\#\Gamma}{\#Y}\geq\aleph$;
  3. (iii)
    #$\#\rho^{-1}(x)$ is the same for all $x\in X$.
The next result corresponds to [BDG10, Lem. 2.5], which is an improved version of [DG09, Prop. 2.5]. Compare also with [${}^{+}$05, Lem. 2 and Section 8].
Theorem 2.3.
Consider a sequence $\{X_{n}\}_{n\in\mathbb{N}}$ of periodic orbits $X_{n}$ of $G$ with increasing periods $P_{n}$. Let $\mu_{n}$ be the uniformly distributed atomic measure over the orbit $X_{n}$. Suppose there exist sequences of real numbers $\{\gamma_{n}\}_{n\in\mathbb{N}}$, $\gamma_{n}>0$, and $\{\aleph_{n}\}_{n\in\mathbb{N}}$, $\aleph_{n}\in(0,1]$, such that:
  1. (1)
    For each $n\in\mathbb{N}$, the orbit $X_{n+1}$ is a $(\gamma_{n},\aleph_{n})$-good approximation of $X_{n}$;
  2. (2)
    $\sum_{n=1}^{\infty}{\gamma_{n}}<\infty$;
  3. (3)
    $\prod_{n=1}^{\infty}{\aleph_{n}}\in(0,1]$.
Then the sequence $\{\mu_{n}\}_{n\in\mathbb{N}}$ has a limit $\mu$. The limit measure $\mu$ is ergodic and the topological support of $\mu$ is the topological limit of the sequence of orbits $\{X_{n}\}_{n\in\mathbb{N}}$, i.e.,
$\mathrm{supp}~{}\mu=\bigcap_{k=1}^{\infty}{\left(\overline{\bigcup_{l=k}^{\infty}{X_{l}}}\right)}.$
One advantage of using an atomic periodic measure for the skew-product $F=\sigma\ltimes\mathscr{F}$ is that it is straightforward to compute its Lyapunov exponent. As a result, we have the following proposition:
Proposition 2.4.
Consider $f\in\langle\mathscr{F}\rangle^{+}$ with a fixed point $x\in\mathbb{S}^{1}$. Then $f$ generates a $P$-periodic orbit $X$ of the skew-product $F=\sigma\ltimes\mathscr{F}$, and
$\lambda(\mu_{X})=\frac{1}{P}\log|f^{\prime}(x)|$
where $\mu_{X}$ denotes the uniformly distributed atomic measure over $X$ and $P$ is the number of maps in $\mathscr{F}$ composing $f$.
We also need to prove Theorem B some others results on properties of an IFS of circle diffeomorphisms collected from the literature. To provide precise statements, we introduce first some notation. Given $\bar{\omega}=\bar{\omega}_{0}\bar{\omega}_{1}\dots\bar{\omega}_{P-1}$ a finite word in the alphabet $\{1,\dots,k\}$, i.e., a sequence of length $|\bar{\omega}|=P$, where $\bar{\omega}_{j}\in\{1,\dots,k\}$ for $j=0,1,\dots,P-1$, we define the cylinder $\llbracket\bar{\omega}\rrbracket\stackrel{{\scriptstyle\scriptscriptstyle\rm def}}{{=}}\{\omega=(\omega_{i})_{i\geq 0}\in\Sigma_{k}~{}:~{}\omega_{i}=\bar{\omega}_{i},~{}~{}i=0,\dots,|\bar{\omega}|-1\}$, and we denote
$T_{\bar{\omega}}\stackrel{{\scriptstyle\scriptscriptstyle\rm def}}{{=}}f_{\bar{\omega}_{P-1}}\circ f_{\bar{\omega}_{P-2}}\circ\dots\circ f_{\bar{\omega}_{0}}.$
The following lemma was proved in [BBD13, Lem. 6.3].
Lemma 2.5.
Assume that $\mathscr{F}$ is minimal. Then for every $\eta>0$ and every non-empty open set $J\subset\mathbb{S}^{1}$, there exist $\delta_{0}=\delta_{0}(\eta,J)>0$ and $K=K(\eta,J)\in\mathbb{N}$ such that for every open interval $I\subset\mathbb{S}^{1}$ of radius less or equal $\delta_{0}$, there exists a finite word $\bar{\omega}$ in the alphabet $\{1,2,\dots,k\}$ with $|\bar{\omega}|=R\leq K$ such that
  1. (I)
    For every $(\omega,x)\in\llbracket\bar{\omega}\rrbracket\times I$, the segment of the orbit of the $F=\sigma\ltimes\mathscr{F}\in S^{r}_{k}(\mathbb{S}^{1})$,
    $F^{[0,R]}(\omega,x)\stackrel{{\scriptstyle\scriptscriptstyle\rm def}}{{=}}\{F^{j}(\omega,x)~{}:~{}0\leq j\leq R\}$
    is $\eta$-dense in $\Sigma_{k}\times\mathbb{S}^{1}$;
  2. (II)
    $T_{\bar{\omega}}(I)\subset J$.
The following Proposition was stated in [${}^{+}$05, Prop. 3].
Proposition 2.6.
If $\mathscr{F}$ is minimal, then for each interval $J$ in $\mathbb{S}^{1}$, there exist $K_{0}=K_{0}(J)\in\mathbb{N}$ and $\delta_{0}=\delta_{0}(J)>0$ such that any interval $I$ of length smaller than $\delta_{0}$ can be brought inside $J$ by some composition of at least $K_{0}$ iterates of the generators of $\mathscr{F}$.
Finally, the next proposition follows immediately from [${}^{+}$17, Lem. 3.1].
Proposition 2.7.
Assume that $\mathscr{F}$ is backward expanding. Then there exist $\nu>1$, $h_{1},\dots,h_{m}\in\langle\mathscr{F}\rangle^{+}$, and $B_{1},\dots,B_{m}$ open intervals of $\mathbb{S}^{1}$ such that
  1. (a)
    $\mathbb{S}^{1}=B_{1}\cup B_{2}\cup\dots\cup B_{m}$.
  2. (b)
    $|h^{\prime}_{i}(x)|>\nu$ for all $x\in B_{i}$ and $i=1,2,\dots,m$.
2.2. Main Lemma
To understand the next lemma, recall the definition of a good approximation given in the previous subsection (Definition 2.2). This lemma is an improved version of [${}^{+}$05, Lemma 3] where we spread the constructed periodic orbits densely on $\Sigma_{k}\times\mathbb{S}^{1}$.
Lemma 2.8.
Consider $F=\sigma\ltimes\mathscr{F}$ and assume that
  1. (i)
    $\mathscr{F}$ is minimal ;
  2. (ii)
    $\mathscr{F}$ is backward expanding;
  3. (iii)
    there exists $f\in\langle\mathscr{F}\rangle^{+}$ with a hyperbolic attracting periodic point.
Then, there exist $10<c<$ 1 and $d>0$ depending only on $\mathscr{F}$, such that if $X$ is the $P$-periodic orbit of $F$ generated by $f$, then for every $\gamma>0$ and $\eta>0$, there exists a periodic orbit $X^{\prime}$ of $F$ with period $P^{\prime}>P$ with the following properties:
  1. (1)
    $c\cdot\lambda(\mu_{X})<\lambda(\mu_{X^{\prime}})<0$;
  2. (2)
    $X^{\prime}$ is $(\gamma,\aleph)$-good approximation [^2] Note that a priori, $\aleph$ may not be positive (requirement of the definition of good approximation). However, the lemma provides a new periodic orbit with smaller exponent. Thus, recursively applying the lemma, we always find an element $f$ satisfying (iii) whose associated periodic orbit has an exponent small enough with $\aleph>0$. of $X$ where =$\aleph\stackrel{{\scriptstyle\scriptscriptstyle\rm def}}{{=}}1-d\cdot|\lambda(\mu_{X})|$;
  3. (3)
    $X^{\prime}$ is $\eta$-dense in $\Sigma_{k}\times\mathbb{S}^{1}$.
Proof.
First, let us prove the existence of the periodic orbit $X^{\prime}$ of $F$ with a period greater than $P$. Since $f\in\langle\mathscr{F}\rangle^{+}$, there exists a finite word $\hat{\omega}$ such that $f=T_{\hat{\omega}}$ and $P=|\hat{\omega}|$. Let $x$ be the attracting periodic point for $f$, meaning $f(x)=x$ and $|f^{\prime}(x)|=\alpha$ where $10<\alpha<$ 1. Let us take constants $\alpha_{-}$ and $\alpha_{+}$ such that
$10<\alpha_{-}<\alpha<\alpha_{+}<$ 1 . (2)
Fix $\gamma>0$. By the continuity of $f^{\prime}$, there is an interval $J$ containing $x$ such that
$0<\alpha_{-}\leq|f^{\prime}(y)|\leq\alpha_{+}<1\quad\text{for all $y\in J$},\quad\text{and}\quad L^{2P}|J|<\gamma,$ (3)
where we denote by $|I|$ the length of an interval $I\subset\mathbb{S}^{1}$ and $L\stackrel{{\scriptstyle\scriptscriptstyle\rm def}}{{=}}\max\{\|g^{\prime}\|_{\infty}:g\in\mathscr{F}\}$.
From Hypothesis (i) and Lemma 2.5, for $J$ and $\eta>0$, there exist $\delta_{0}=\delta_{0}(J,\eta)>0$ and $K=K(J,\eta)\in\mathbb{N}$ such that (I) and (II) hold. On the other hand, from Hypothesis (ii) and Proposition 2.7, there exist a constant $\nu>1$, diffeomorphisms $h_{1},\dots,h_{m}\in\langle\mathscr{F}\rangle^{+}$, and open intervals $B_{1},\dots,B_{m}$ of $\mathbb{S}^{1}$ such that (a) and (b) hold. Let $\lambda>0$ be the Lebesgue number of the covering in (a). Recall that this means any interval of length less than $\lambda$ is contained in some $B_{i}$. Let us set
$L_{1}=\max{\|h^{\prime}_{i}\|_{\infty}:i=1,\dots,m\}}>1\quad\text{and}\quad\delta=\min\{\delta_{0},|J|L^{-K},\lambda/L_{1}\}>0.$ (4)
For each $n\in\mathbb{N}$ sufficiently large, we take $r=r(n)\in\mathbb{N}$ to be the unique integer such that
$0<\frac{\log\frac{\alpha_{+}^{-n}\delta}{L_{1}|J|}}{\log L_{1}}\leq r<\frac{\log\frac{\alpha_{+}^{-n}\delta}{L_{1}|J|}}{\log L_{1}}+1=\frac{\log\frac{\alpha_{+}^{-n}\delta}{|J|}}{\log L_{1}}.$ (5)
Note that it holds
$\frac{\delta}{L_{1}}\leq\alpha_{+}^{n}L_{1}^{r}|J|<\delta.$ (6)
Since $x\in J$ is an attracting fixed point of $f$, using the Intermediate Value Theorem and (3), (6) and (4), we have $|f^{n}(J)|\leq\alpha_{+}^{n}|J|\leq\alpha^{n}_{+}L_{1}^{r}|J|<\delta<\lambda$. Thus, there exists $i_{1}\in\{1,\dots,m\}$ such that $f^{n}(J)\subset B_{i_{1}}$. Similarly, since
$|h_{i_{1}}f^{n}(J)|\leq\alpha_{+}^{n}L_{1}|J|\leq\alpha_{+}^{n}L_{1}^{r}|J|<\delta<\lambda,$
we obtain $i_{2}\in\{1,\dots,m\}$ such that $h_{i_{1}}f^{n}(J)\subset B_{i_{2}}$. Proceeding recursively, we obtain $i_{1},\dots,i_{r}\in\{1,\dots,m\}$ such that
$f^{n}(J)\subset B_{i_{1}}\ \ \text{and}\ \ h_{i_{j}}\dots h_{i_{1}}f^{n}(J)\subset B_{i_{j+1}}\ \text{for every $j=1,\dots,r-1$.}$ (7)
[…]
$\alpha^{n}_{-}\nu^{r}|J|\leq|T_{\bar{\kappa}_{1}}T_{\bar{\kappa}_{0}}(J)|\leq L_{1}^{r}\alpha_{+}^{n}|J|.$ (8)
Let $I=T_{\bar{\kappa}_{1}}T_{\bar{\kappa}_{0}}(J)$. Then, from equations (6), (8), it follows that $|I|<\delta\leq\delta_{0}$. Thus, by the definition of $\delta_{0}=\delta_{0}(J,\eta)$ and (II), there exists a finite word $\bar{\kappa}_{2}$, with $|\bar{\kappa}_{2}|\leq K$, such that $T_{\bar{\kappa}_{2}}(I)\subset J$. Therefore, $T_{\bar{\kappa}_{2}}T_{\bar{\kappa}_{1}}(J)\subset J$ by the definition of $I$. Taking the finite word $\bar{\omega}=\bar{\kappa}_{0}\bar{\kappa}_{1}\bar{\kappa}_{2}$, we have $T_{\bar{\omega}}=T_{\bar{\kappa}_{2}}\circ T_{\bar{\kappa}_{1}}\circ T_{\bar{\kappa}_{0}}$, and thus,
$T_{\bar{\omega}}(J)\subset J\quad\text{with}\quad|\bar{\omega}|=nP+|\bar{\kappa}_{1}|+|\bar{\kappa}_{2}|>P.$ (9)
From $|\bar{\kappa}_{2}|\leq K$, (4), and (6), we have
$|T^{\prime}_{\bar{\omega}}(y)|\leq L^{K}L_{1}^{r}\alpha_{+}^{n}<L^{K}\frac{\delta}{|J|}\leq L^{K}\frac{|J|L^{-K}}{|J|}=1\quad\text{for all}\quad y\in J.$ (10)
Hence, (9) and (10) imply that $T_{\bar{\omega}}$ is a contraction on $J$. Therefore, by the Banach Fixed-Point Theorem, $T_{\bar{\omega}}$ has a unique fixed point $x^{\prime}\in J$. Finally, by Proposition 2.4, there exists a periodic orbit $X^{\prime}$ generated by $T_{\bar{\omega}}$ with period $P^{\prime}=|\bar{\omega}|>P$.
It should be noted that the previous argument holds for any sufficiently large integer $n$ and constants $\alpha_{+}$ and $\alpha_{-}$ satisfying (2). Furthermore, since $\bar{\kappa}_{2}$ is the finite word given by Lemma 2.5 and the fact that $\sigma^{nP+\bar{\kappa}_{1}}(\bar{\omega}\bar{\omega}\dots)\in\llbracket\bar{\kappa}_{2}\rrbracket$, it follows from (I) that the tail of the orbit $X^{\prime}$ is $\eta$-dense in $\Sigma_{k}\times\mathbb{S}^{1}$. In particular, the orbit $X^{\prime}$ is $\eta$-dense in $\Sigma_{k}\times\mathbb{S}^{1}$, thus proving Item (3).
Now let us calculate the Lyapunov exponent along the fiber with respect to the measure $\mu_{X^{\prime}}$. For this, according to (8), we have $|T^{\prime}_{\bar{\omega}}(x^{\prime})|\geq|T^{\prime}_{\bar{\kappa}_{2}}(T_{\bar{\kappa}_{1}}T_{\bar{\kappa}_{0}}(x^{\prime}))|\,\alpha_{-}^{n}\nu^{r}\geq e^{C_{1}}\alpha_{-}^{n}\nu^{r}$, where
$C_{1}=\min\{\log|DT_{\bar{\kappa}}(y)|:\,y\in\mathbb{S}^{1}\ \text{and}\ |\bar{\kappa}|<K\}.$ (11)
Thus, we have
$\log|T^{\prime}_{\bar{\omega}}(x^{\prime})|>C_{1}+r\log\nu+n\log\alpha_{-}.$ (12)
Substituting (5) into (12), we get
$\log|T^{\prime}_{\bar{\omega}}(x^{\prime})|>C_{1}+\frac{\log\nu}{\log L_{1}}\log\frac{\alpha_{+}^{-n}\delta}{L_{1}|J|}+n\log\alpha_{-}=C_{1}+C_{2}+n\,(\log\alpha_{-}-\frac{\log\nu}{\log L_{1}}\log\alpha_{+})$ (13)
where
$C_{2}=\frac{\log\nu}{\log L_{1}}\log\frac{\delta}{L_{1}|J|}.$ (14)
Taking into account that $10<\alpha<$ 1 and $1<\nu<L_{1}$,
$(1-\frac{2\log\nu}{3\log L_{1}})\log\alpha<(1-\frac{\log\nu}{\log L_{1}})\log\alpha<0.$
Thus, since
$\lim_{(\alpha_{-},\alpha_{+})\to(\alpha,\alpha)}\log\alpha_{-}-\frac{\log\nu}{\log L_{1}}\log\alpha_{+}=(1-\frac{\log\nu}{\log L_{1}})\log\alpha,$
there exist $\alpha_{-}$ and $\alpha_{+}$ satisfying (2) and (3), such that
$(1-\frac{2\log\nu}{3\log L_{1}})\log\alpha\leq\log\alpha_{-}-\frac{\log\nu}{\log L_{1}}\log\alpha_{+}.$ (15)
Writing $C_{3}=C_{1}+C_{2}$ and substituting (15) into (13), it follows that
$\log|T^{\prime}_{\bar{\omega}}(x^{\prime})|>C_{3}+n(1-\frac{2\log\nu}{3\log L_{1}})\log\alpha.$ (16)
From Proposition 2.4, the Lyapunov exponent along the fiber with respect to the measure $\mu_{X^{\prime}}$ can be estimated as
$\lambda(\mu_{X^{\prime}})=\frac{1}{P^{\prime}}\log|T^{\prime}_{\bar{\omega}}(x^{\prime})|=\frac{\log|T^{\prime}_{\bar{\omega}}(x^{\prime})|}{nP+|\bar{\kappa}_{1}|+|\bar{\kappa}_{2}|}.$
Noting that, $\log|T^{\prime}_{\bar{\omega}}(x^{\prime})|<0$ since $T_{\bar{\omega}}$ is a contraction, and substituting (16) into this last equation, we have
$\lambda(\mu_{X^{\prime}})>\frac{1}{nP}\left(C_{3}+n(1-\frac{2\log\nu}{3\log L_{1}})\log\alpha\right)=\lambda(\mu_{X})(1-\frac{2\log\nu}{3\log L_{1}})+\frac{C_{3}}{nP}$ (17)
where the last equality follows since $\lambda(\mu_{X})=(\log\alpha)/P$. Since $C_{3}=C_{2}+C_{1}$ and the constants $C_{1}$ and $C_{2}$ in (11) and (14) do not depend on $n$ (they only depend on $\mathscr{F}$ and the interval $J$), then $C_{3}/nP\to 0$ as $n\to\infty$. Thus, similarly as above, since
$\lambda(\mu_{X})(1-\frac{\log\nu}{2\log L_{1}})<\lambda(\mu_{X})(1-\frac{2\log\nu}{3\log L_{1}}),$
we find a sufficiently large $n$ such that (17) becomes
$\lambda(\mu_{X^{\prime}})>\lambda(\mu_{X})(1-\frac{2\log\nu}{3\log L_{1}})+\frac{C_{3}}{nP}>\lambda(\mu_{X})(1-\frac{\log\nu}{2\log L_{1}}).$
This proves Item (1) of the lemma with $c=1-\frac{\log\nu}{2\log L_{1}}\in(0,1)$.
[…]
$\tilde{X^{\prime}}=\{F^{j}(\omega^{\prime},x^{\prime}):\,MP\leq j<(n-M-1)P\},$ (18)
is contained in the set of the first $n$ iterations by $F^{P}$ of the point $(\omega^{\prime},x^{\prime})$. Now let us define the projection $\rho:\tilde{X^{\prime}}\to X$ as follows:
$\rho(F^{j}(\omega^{\prime},x^{\prime}))=F^{j}(\omega,x).$ (19)
Then, for $i=0,\dots,P-1$, $j=MP,\dots,(n-M-1)P-1$, we have
$d\left(F^{i}\left(F^{j}(\omega^{\prime},x^{\prime})\right),F^{i}\left(\rho\circ F^{j}(\omega^{\prime},x^{\prime})\right)\right)$ $=d\left(F^{i+j}(\omega^{\prime},x^{\prime}),F^{i+j}(\omega,x)\right)$
$=\max\{\bar{d}(\sigma^{i+j}(\omega^{\prime}),\sigma^{i+j}(\omega)),|f^{i+j}_{\omega^{\prime}}(x^{\prime})-f^{i+j}_{\omega}(x)|\}.$
Since the first $nP$ letters of $\omega^{\prime}$ and $\omega$ are the same, and since $MP\leq i+j<nP$, we have $\bar{d}(\sigma^{i+j}(\omega^{\prime}),\sigma^{i+j}(\omega))\leq 2^{-MP}<\gamma$. For the same reason, writing $j=qP+\tau$, we have
$f^{i+j}_{\omega^{\prime}}=f^{i+j}_{\omega}=f^{i}_{\sigma^{j}(\omega)}\circ f^{j}_{\omega}=f^{i}_{\sigma^{\tau}(\omega)}\circ f^{\tau}_{\omega}\circ f^{q}=f^{i+\tau}_{\omega}\circ f^{q}.$
Then, since $x$ and $x^{\prime}$ are in $J$ and $f^{q}(J)\subset J$, by virtue of (3), it holds tha $|f^{i+j}_{\omega^{\prime}}(x^{\prime})-f^{i+j}_{\omega}(x)|\leq|f^{i+\tau}_{\omega}(J)|\leq L^{2P}|J|<\gamma$. This concludes that
$d\left(F^{i}(y),F^{i}(\rho(y))\right)<\gamma\quad\text{for all $y\in\tilde{X^{\prime}}$ and $i=0,\dots,P-1$}.$
proving Item (i) in Definition 2.2.
Now let us show Item (ii) in Definition 2.2. For this, note that
$1-\frac{\#\tilde{X^{\prime}}}{\#X^{\prime}}$ $=\frac{|\bar{\kappa}_{1}|+|\bar{\kappa}_{2}|+(2M+1)P}{nP+|\bar{\kappa}_{1}|+|\bar{\kappa}_{2}|}\leq\frac{|\bar{\kappa}_{1}|+K+(2M+1)P}{nP}.$ (20)
On the other hand, denoting by $H=\max\{|\bar{\beta}_{1}|,\dots,|\bar{\beta}_{m}|\}$ where $\bar{\beta}_{i}$ is the finite word that generates $h_{i}$, by virtue of (4), (5), and $\bar{\kappa}_{1}=\bar{\beta}_{i_{r}}\dots\bar{\beta}_{i_{1}}$, we have
$|\bar{\kappa}_{1}|$ $\leq rH\leq\frac{-n\log\alpha_{+}+\log\delta|J|^{-1}}{\log L_{1}}H$
$\leq\left(-n\log\alpha-K\log L\right)\frac{H}{\log L_{1}}=-\left(n{P\,\lambda(\mu_{X})}+{K\log L}\right)\frac{H}{\log L_{1}}.$
Substituting this expression into (20) we have
$1-\frac{\#\tilde{X^{\prime}}}{\#X^{\prime}}\leq|\lambda(\mu_{X})|\frac{H}{\log L_{1}}+\frac{C_{4}}{nP}~{},$
where $C_{4}=K+(2M+1)P-KH\frac{\log L}{\log L_{1}}$. Since $C_{4}$ does not depend on $n$, by again taking this integer sufficiently large, it follows that
$\frac{C_{4}}{nP}+|\lambda(\mu_{X})|\frac{H}{\log L_{1}}\leq|\lambda(\mu_{X})|\frac{2H}{\log L_{1}}.$
Therefore, from (20), we have
$\aleph=1-d\cdot|\lambda(\mu_{X})|\leq\frac{\#\tilde{X^{\prime}}}{\#X^{\prime}}\quad\text{where $d=2H/\log L_{1}>0$}.$
This shows Item (ii) in Definition 2.2.
Finally, we only need to show Item (iii) in Definition 2.2. For this, given $\tau=0,1,\dots,P-1$, we have
#$\#\rho^{-1}(F^{\tau}(\omega,x))=$$\#\{F^{j}(\omega^{\prime},x^{\prime})\in\tilde{X}^{\prime}:\,\rho(F^{j}(\omega^{\prime},x^{\prime}))=F^{\tau}(\omega,x)\}$
$=$ #$\#\{j\in\{MP\leq j<(n-M-1)P\}:\,j=\mathring{P}+\tau\}$
$=$ $\frac{(n-M-1)P-MP}{P}=n-2M-1.$
This proves the last item and therefore $X^{\prime}$ is a $(\gamma,\aleph)$-good approximation of $X$. ∎
2.3. Proof
Finally, we will prove the main result of this section.
Proof of Theorem B.
To prove this result, we will use Theorem 2.3. First, let us show that $F$ has a sequence of periodic orbits $\{X_{n}\}$ with increasing period $P_{n}$.
Let $\{\gamma_{n}\}_{n\in\mathbb{N}}$ and $\{\eta_{n}\}_{n\in\mathbb{N}}$ be positive sequences such that
$\sum_{n=1}^{\infty}{\gamma_{n}}<\infty~{}~{}\text{and}~{}~{}\eta_{n}\to 0~{}~{}\text{as}~{}~{}n\to\infty.$ (21)
Let $X_{1}$ be the periodic orbit of $F$ generated by $f$ with period $P_{1}$. Recursively applying Lemma 2.8, for each $\gamma_{n}>0$ and $\eta_{n}>0$, there exists a periodic orbit $X_{n+1}$ of $F$ with period $P_{n+1}>P_{n}$ such that
  1. a)
    $c\cdot\lambda(\mu_{X_{n}})<\lambda(\mu_{X_{n+1}})<0$;
  2. b)
    $X_{n+1}$ is a $(\gamma_{n},\aleph_{n})$-Good approximation of $X_{n}$ with $\aleph_{n}=1-d\cdot|\lambda(\mu_{X_{n}})|$;
  3. c)
    $X_{n+1}$ is $\eta_{n}$-dense in $\Sigma_{k}\times\mathbb{S}^{1}$.
where $10<c<$ 1 and $d>0$ are constants depending only on $\mathscr{F}$. In particular, $c^{n}\cdot\lambda(\mu_{X_{1}})<\lambda(\mu_{X_{n+1}})<0$. Thus, $\lambda(\mu_{X_{n+1}})\to 0$ as $n\to\infty$, and therefore $\aleph_{n}>0$ for sufficiently large $n$. Without loss of generality, we assume that this holds for all $n\geq 1$. Furthermore, since
$\sum_{n=1}^{\infty}|\lambda(\mu_{X_{n}})|\leq\sum_{n=0}^{\infty}c^{n}|\lambda(\mu_{X_{1}})|<\infty,$
by the Comparison Test for Series, we also have
$-\infty<\sum_{n=1}^{\infty}\log\left(1-d|\lambda(\mu_{X_{n}})|\right)=\sum_{n=1}^{\infty}\log\aleph_{n}<0,$
and consequently $\prod_{n=1}^{\infty}{\aleph_{n}}\in(0,1]$. By Theorem 2.3, the sequence $\{\mu_{X_{n}}\}$ of atomic measures uniformly distributed over the orbit $X_{n}$ has a limit, and the limit measure $\mu$ is ergodic with $\mathrm{supp}~{}\mu=\bigcap_{j=1}^{\infty}{\left(\overline{\bigcup_{l=j}^{\infty}{X_{l}}}\right)}$. From c) and the fact that $\eta_{n}\to 0$ as $n\to\infty$, we have $\overline{\bigcup_{l=j}^{\infty}{X_{l}}}=\Sigma_{k}\times\mathbb{S}^{1}$, which implies $\mathrm{supp}~{}\mu=\Sigma_{k}\times\mathbb{S}^{1}$.
On the other hand, since the sequence of Lyapunov exponents $\{\lambda(\mu_{X_{n}})\}$ converges to zero, it follows from Proposition 2.1 that $\lambda(\mu)=0$. This completes the proof of the theorem. ∎
3. Proof of Theorem A
The following result corresponds to [BC23, Thm. 3.22], which improves [BFS14, Thm. B]. This generalizes earlier results on robustly minimal IFSs on the circle described in [GI99, GI00] (see also [KKO18, Pro. 6.1]). Recall that an IFS is said to be robustly minimal (transitive) if any sufficiently small $C^{1}$-perturbation remains minimal (transitive).
Theorem 3.1.
Let $f_{1},f_{2}$ be preserving orientation circle diffeomorphisms. Assume that $f_{1}$ is an irrational rotation and that $f_{2}$ is not a rotation. Then =$\mathscr{F}=\{f_{1},f_{2}\}$ is, both, forward and backward expanding and robust forward and backward minimal.
The next result was proved in [BC23, Cor. 5.5] as a consequence of previous results by Kleptsyn, Kudryashov and Okunev in [KKO18, Thm. 1.5, Cor. 5.7]. First, we recall that a family $\mathscr{F}$ in $\mathrm{IFS}^{r}_{k}(\mathbb{S}^{1})$ is transitive if, and only if, the associated skew-product $F=\sigma\ltimes\mathscr{F}$ is transitive. For equivalent definitions see [BC23, Def. 4.1, Thm. 4.2, 4.5 and 6.1].
Theorem 3.2.
Let $\mathscr{F}$ be a transitive family in $\operatorname{\mathrm{IFS}}^{r}_{k}(\mathbb{S}^{1})$, where $r\geq 1$ and $k\geq 2$. Then, there exists $\mathscr{G}\in\operatorname{\mathrm{IFS}}^{r}_{k}(\mathbb{S}^{1})$, $C^{r}$ close to $\mathscr{F}$, such that $\mathscr{G}$ is both robustly forward and backward minimal, the elements of $\mathscr{G}$ are Morse-Smale diffeomorphisms, and there is a map $g_{0}\in\langle\mathscr{G}\rangle^{+}$ that is $C^{\infty}$ conjugate to an irrational rotation.
Now we are ready to give the proof of our main result.
Proof of Theorem A.
Denote by $\mathrm{T}^{r}_{k}(\mathbb{S}^{1})$ the set of transitive family in $\mathrm{IFS}^{r}_{k}(\mathbb{S}^{1})$. Since every minimal IFS is also transitive (cf. [BC23, Rem. 4.4]), by Theorem 3.2, there is a dense set $\mathcal{D}$ of $\mathrm{T}^{r}_{k}(\mathbb{S}^{1})$ such that every $\mathscr{G}\in\mathcal{D}$ is robustly transitive, its elements are Morse-Smale diffeomorphisms, and there is a map $g_{0}\in\langle\mathscr{G}\rangle^{+}$ that is $C^{\infty}$ conjugate to an irrational rotation. Let $\phi$ be the $C^{\infty}$ conjugacy map between $g_{0}$ and the rigid irrational rotation. Set $\hat{\mathscr{G}}=\phi\circ\mathscr{G}\circ\phi^{-1}$. Note that since the property of being a Morse-Smale diffeomorphism is invariant under smooth conjugation, the elements of $\hat{\mathscr{G}}$ are also Morse-Smale diffeomorphisms. Moreover, since $\langle\hat{\mathscr{G}}\rangle^{+}$ also has an irrational rotation, by Theorem 3.1, we conclude that $\hat{\mathscr{G}}$ is both backward and forward expanding and robustly forward and backward minimal. Since the expanding property is also a robust property (cf. [${}^{+}$17, Cor. 3.1] or [BC23, Rem. 3.10]) and $\mathscr{G}$ is robustly transitive, we can consider a neighborhood $\mathcal{U}_{\mathscr{G}}\subset\mathrm{T}^{r}_{k}(\mathbb{S}^{1})$ of $\mathscr{G}$ in $\operatorname{\mathrm{IFS}}^{r}_{k}(\mathbb{S}^{1})$ such that for every $\mathscr{H}\in\mathcal{U}_{\mathscr{G}}$, the family ^$\hat{\mathscr{H}}=\phi\circ\mathscr{H}\circ\phi^{-1}$ is both backward and forward expanding and robustly forward and backward minimal. Thus, by Theorem B, the skew-product $\hat{H}=\sigma\ltimes\hat{\mathscr{H}}$ has a $\hat{H}$-invariant ergodic probability measure $\hat{\mu}$ with full support, limit of periodic measures and with Lyapunov exponent $\lambda(\hat{\mu})=0$. Consider the map $\Phi=\mathrm{id}_{\Sigma_{k}}\times\phi$ and notice that $\hat{H}=\Phi\circ H\circ\Phi^{-1}$, where $H=\sigma\ltimes\mathscr{H}$.
Claim 3.3.
The measure $\mu\stackrel{{\scriptstyle\scriptscriptstyle\rm def}}{{=}}\hat{\mu}\circ\Phi$ is a $H$-invariant ergodic probability with full support, limit of periodic measures and with Lyapunov exponent $\lambda(\mu)=0$.
Proof.
We first observe that $\hat{H}^{-1}=\Phi\circ H^{-1}\circ\Phi^{-1}$. Hence, for any measurable set $A\subset\Sigma_{k}\times\mathbb{S}^{1}$, since $\hat{\mu}$ is a $\hat{H}$-invariant measure, i.e., $(\hat{H})_{*}\hat{\mu}=\hat{\mu}$, we have
$\mu(A)=\hat{\mu}(\Phi(A))=(\hat{H})_{*}\hat{\mu}(\Phi(A))=\hat{\mu}(\Phi\circ H^{-1}\circ\Phi^{-1}\circ\Phi(A))=\hat{\mu}(\Phi\circ H^{-1}(A))=\mu(H^{-1}(A)).$
Thus, $\mu$ is $H$-invariant. Moreover, if $A$ satisfies $H^{-1}(A)=A$, then
$\hat{H}^{-1}(\Phi(A))=\Phi\circ H^{-1}\circ\Phi^{-1}(\Phi(A))=\Phi\circ H^{-1}(A)=\Phi(A)$
and thus, since $\hat{\mu}$ is an ergodic measure for $\hat{H}$, we have that $\mu(A)=\hat{\mu}(\Phi(A))\in\{0,1\}$. Hence, $\mu$ is also an ergodic measure for $H$. Also, since $\Phi$ is a homeomorphism and $\hat{\mu}$ has full support, then $\mu=\hat{\mu}\circ\Phi$ inherits full support. Furthermore, $\mu$ is also a limit of periodic measures, as $\hat{\mu}$ is. Finally, we will calculate the Lyapunov exponent $\lambda(\mu)$. To do this, observe that ^$\hat{\mathscr{H}}=\phi\circ\mathscr{H}\circ\phi^{-1}$ and thus the fiber maps of ^$\hat{\mathscr{H}}^{n}$ are given by $\hat{h}^{n}_{\omega}=\phi\circ h^{n}_{\omega}\circ\phi^{-1}$ for any $\omega\in\Sigma_{k}$, where $h^{n}_{\omega}$ are the fiber maps of $H^{n}$. Hence,
$\lambda_{\hat{H}}(\omega,x)$ $=\lim_{n\to\infty}\frac{1}{n}\log|(\hat{h}^{n}_{\omega})^{\prime}(x)|$
$=\lim_{n\to\infty}\frac{1}{n}\log|\phi^{\prime}(h^{n}_{\omega}\circ\phi^{-1}(x))|+\lim_{n\to\infty}\frac{1}{n}\log|(h^{n}_{\omega})^{\prime}(\phi^{-1}(x))|+\lim_{n\to\infty}\frac{1}{n}\log|(\phi^{-1})^{\prime}(x)|$
$=\lim_{n\to\infty}\frac{1}{n}\log|(h^{n}_{\omega})^{\prime}(\phi^{-1}(x))|=\lambda_{H}(\omega,\phi^{-1}(x))=\lambda_{H}(\Phi^{-1}(\omega,x))$
for every $(\omega,x)\in\Sigma_{k}\times\mathbb{S}^{1}$. Now, since $\hat{\mu}=\mu\circ\Phi^{-1}$, by the above equality, we have
[…]
Since $\lambda(\hat{\mu})=0$, we also get that $\lambda(\mu)=0$ as required. This completes the proof ∎
Now, we conclude the proof by taking $\tilde{\mathcal{R}}$ as the open set formed by the union of the neighborhoods $\mathcal{U}_{\mathscr{G}}$ for $\mathscr{G}\in\mathcal{D}$. Since $\mathcal{U}_{\mathscr{G}}\subset\mathrm{T}^{r}_{k}(\mathbb{S}^{1})$ and $\mathcal{D}$ is dense in $\mathrm{T}^{r}_{k}(\mathbb{S}^{1})$, we get that $\tilde{\mathcal{R}}$ is open and dense in $\mathrm{T}^{r}_{k}(\mathbb{S}^{1})$. Then, by the definition of the topology in $\mathrm{S}^{r}_{k}(\mathbb{S}^{1})$ and in view that $\mathscr{F}\in\mathrm{T}^{r}_{k}(\mathbb{S}^{1})$ if and only if $F=\sigma\ltimes\mathscr{F}\in\mathrm{T}^{r}_{k}(\mathbb{S}^{1})$, we get the set =$\mathcal{R}=\{F=\sigma\ltimes\mathscr{F}\in\mathrm{S}^{r}_{k}(\mathbb{S}^{1}):\mathscr{F}\in\tilde{\mathcal{R}}\}$ is an open and dense subset of $\mathrm{TS}^{r}_{k}(\mathbb{S}^{1})$. Moreover, for every $H=\sigma\ltimes\mathscr{H}\in\mathcal{R}$, the family $\mathscr{H}$ belongs to $\mathcal{U}_{\mathscr{G}}$ for some $\mathscr{G}\in\mathcal{D}$ and thus by Claim 3.3, there is an ergodic $F$-invariant probability measure $\mu$ with full support whose Lyapunov exponent $\lambda(\mu)$ equal to zero. Furthermore, the probability measure $\mu$ is a limit in the $*{}^{*}$ topology of sequences of invariant measures supported on periodic points. This concludes the proof. ∎
Acknowledgement
The proof of Theorem B forms part of the results obtained in the master’s thesis of the author, J. A. Cisneros. The first author, P. G. Barrientos, was supported by grant PID2020-113052GB-I00 funded by MCIN, PQ 305352/2020-2 (CNPq), and JCNE E-26/201.305/2022 (FAPERJ). We would like to express our deepest gratitude to Prof. Lorenzo J. Díaz for his insightful comments on the results obtained. His contributions as a member of the master’s and doctoral committees of the first author have been invaluable and greatly appreciated.